# Benchmark of segmentation techniques for pelvic fracture in CT and X-ray: Summary of the PENGWIN 2024 challenge

### v1

@@ 4.2.1 CT segmentation
The authors formulated the task as two semantic segmentation problems. The first network performed anatomical segmentation of the three bone regions. The second network predicted the adaptive border boundary core (ABBC) representation, which consists of four semantic classes: background, boundary, border, and core. The core and boundary classes together define individual fragments, with the boundary class acting as a separator between the core and the background or adjacent fragments. The authors introduced a novel dynamic boundary thickness adjustment using the medial axis transform, ensuring that thin fragments remain connected within the core region. The border class was defined as the intersection of the dilated outlines of neighboring fragments. Given a prediction, connected component analysis was applied to the core class to extract initial fragment instances. The boundary class was then merged with the nearest seed to form fragment instances. To ensure anatomical consistency, fragment instances were adjusted by splitting them along anatomical borders and merging smaller fragments with the closest anatomically correct instance. Additionally, the border class was used to refine the segmentation by modifying fragments through split-merge operations, ensuring alignment with the network predictions. [^5] MIC-DKFZ CT algorithm: https://codebase.helmholtz.cloud/hi-dkfz/applied-computer-vision-lab/challenges/abbc
@@
The authors proposed a two-stage segmentation framework. The first stage focused on localizing the three bone regions. To achieve this, the authors trained a U-Net model on downsampled CT scans, which reduced the computational costs and provided a global context that prevented excessive false positives when processing smaller patches. The localized regions were then refined at full resolution for accurate anatomical segmentation. The second stage focused on fracture segmentation within each of the three bone regions. The authors categorized bone fragments into primary (largest fragment) and secondary (smaller fragments) components. After obtaining initial predictions, connected component analysis was applied to reassign secondary fragment labels in accordance with evaluation rules. The final segmentation result was generated by combining the outputs from both stages. A final refinement step ensured that the fracture segmentation was aligned with the foreground mask from the first stage. [^6] SMILE CT algorithm: https://github.com/yuepeiyan/PENGWIN_Challenge
@@
The authors employed a two-stage segmentation pipeline to accurately localize and segment fractured bone fragments. In the first stage, the TotalSegmentator (Wasserthal et al., 2023) model was used to localize the pelvic region, after which the three bone regions were cropped from the original CT scans. In the second stage, the NexToU model was applied to segment primary and secondary fragments within each cropped bone region. The largest connected component was extracted as the primary fragment, and the remaining secondary fragments were expected to be naturally separated. A connected component analysis step was used to isolate individual secondary fragments. [^7] MedIG CT algorithm: https://github.com/pzhhhhh2263/MICCAI-Challenge-PENGWIN2024
@@
The authors developed a refinement-based segmentation pipeline, integrating anatomical segmentation with fracture surface segmentation to ensure anatomically consistent fracture identification. The approach utilized two deep networks. The first network, adapted from the FracSegNet study (Liu et al., 2023), was responsible for segmenting the three bone regions. The second network, trained on fracture surfaces extracted from the PENGWIN dataset, identified fracture boundaries. For training, the authors generated a complete anatomical segmentation of the pelvis by merging all bone fragments. They then extracted contours of both individual fragments and the anatomical segmentation, removing overlapping edges to isolate fracture surfaces—where fractured bones were in contact. The fracture network was trained to predict these fracture surfaces from CT scans. In post-processing, segmentations from both networks were merged. A continuity analysis was then performed to separate individual fractures, with fragments labeled according to their anatomical overlap. [^8] MedApp-AGH CT algorithm: https://github.com/Jarartur/pengwin-challenge-submission
@@
All five teams employed a two-stage approach, first performing anatomical segmentation before segmenting fracture fragments. The key variations among their methods lay in the fracture segmentation step—whether to extract individual fragments directly or to first detect border regions and infer fragment separation. Most teams adopted the nnU-Net framework, owing to its robustness and user-friendly data processing pipeline. Additionally, Dice and cross-entropy (CE) loss were commonly used, as these loss functions have demonstrated effectiveness in various medical imaging tasks.
In Table 3, we summarize some of the main components of these algorithms, including network architecture, pre-processing, and training schemes. A detailed discussion of methodology is included in Sec. 5.1.
4.2.2 X-ray segmentation
Table 4: Summary of the top-performing participating algorithms for X-ray segmentation. Spatial augmentation includes random scaling, mirroring, and flipping. Intensity augmentation includes noising, blurring, and histogram adjustments.
@@ 4.2.2 X-ray segmentation
The authors addressed the limitations of semantic segmentation in handling overlapping anatomical structures in X-ray images by developing a two-stage hierarchical segmentation strategy similar to their CT solution. In the first stage, three dedicated U-Net models were trained to segment the three bone regions separately. The isolated bone regions allowed for focused substructure analysis within each area. In the second stage, a hierarchical fragment classification approach was implemented, adapted from the authors’ CT segmentation method, with modifications specific to X-ray imaging. Separate models were employed to distinguish primary and secondary fracture fragments, mitigating the overlapping structures in X-rays. While this approach effectively reduced primary-secondary overlap, challenges remained in cases where multiple secondary fragments overlapped. For post-processing, connected component analysis was applied to relabel secondary fragments, while primary fragment predictions were refined through consistency checks with the localization masks from the first stage, ensuring that only validated foreground pixels were retained. [^9] SMILE X-ray algorithm: https://github.com/yuepeiyan/PENGWIN_Challenge
@@
The authors employed a multi-model approach to process each of the three bone regions in X-ray images. Each model processed a 2D X-ray image and generated a multi-channel 3D tensor as output. The first channel represents the primary fragment, while the remaining nine channels corresponds to secondary fragments. The segmentation results from the three models were then combined to construct the final output. To optimize model performance, the authors applied a loss function composed of an unweighted combination of Dice loss, binary CE loss, and focal Tversky loss. A pre-trained ImageNet model was used to initialize the network, enhancing feature extraction capabilities. The network architecture was based on a UNet++ backbone with a RegNetY-160 encoder, which provided a balance between spatial resolution and computational efficiency. [^10] MedIG X-ray algorithm: https://github.com/pzhhhhh2263/MICCAI-Challenge-PENGWIN2024
@@
The authors proposed a category and fragment segmentation (CFS) framework. First, three separate category segmentation networks, each based on the Swin-UNETR architecture, were used to segment the three bone regions from X-ray images. After obtaining the anatomical labels, a fragment segmentation network was applied to isolate the bone fragments within each region. The fragment segmentation step employed three separate Mask R-CNN-based networks, which performed both segmentation and object detection through bounding box prediction and classification. These networks took both the X-ray images and the binary anatomical masks as input, allowing the model to focus on specific bone regions while reducing interference from surrounding structures. During inference, the network predicted multiple bounding boxes, each associated with a confidence score. Only bounding boxes with a confidence score above 0.8 were retained for further analysis. In post-processing, the fragment segmentation masks were multiplied by the anatomical masks from the first step to ensure spatial consistency. [^11] LME X-ray algorithm: https://github.com/DaE-plz/CFSSegNet
@@
The authors employed Mask R-CNN with a ResNet50 backbone network for instance segmentation. The model consists of two main stages: a region proposal network that generates candidate object regions, and a detection network that refines these regions while performing anatomical classification, bounding-box regression, and mask prediction. Mask R-CNN was employed to predict pixel-wise binary masks for each region of interest (RoI). To maintain spatial alignment between RoIs and extracted features, the network replaced RoI Pooling with RoI Align, which eliminates quantization errors by using bilinear interpolation instead of discrete binning. This significantly improved mask prediction accuracy, particularly for small objects. To address the specific challenges of pelvic fragment segmentation, the loss weights between classification and segmentation were carefully adjusted during training. Additionally, to improve segmentation accuracy, a structured loss function combining IoU and cross-entropy loss was incorporated, enhancing the model’s robustness and segmentation quality in occluded scenarios.
@@
The solutions in the X-ray task exhibited greater diversity in their overall approach compared to the CT task. The main components of these algorithms are summarized in Table 4 and further discussed in Sec. 5.1.
4.3 Performance analysis
4.3.1 CT segmentation performance
Figure 3: Example CT segmentation results from the top-performing teams. Four representative cases are shown in (a)-(d), illustrating varying levels of complexity from easy to difficult.
Example CT segmentation results are shown in Fig. 3. The performance metrics of each algorithm submitted to the final test are summarized in Table 5 and Fig. 5.
Table 5: Quantitative results for the final test phase of the CT task. HD95 and ASSD values are reported in millimeters. The best scores for each metric are shown in bold.
Team Rank ↓ IoU-F ↑ HD95-F ↓ ASSD-F ↓ IoU-A ↑ HD95-A ↓ ASSD-A ↓
MIC-DKFZ 1 0.9296 5.866 1.843 0.9810 2.368 1.290
SMILE 2 0.9096 5.562 1.663 0.9802 2.410 1.321
MedIG 3 0.9084 6.207 1.680 0.9798 2.452 1.298
MedApp-AGH 4 0.9049 6.899 2.007 0.9783 2.398 1.385
Sano 5 0.8802 6.866 1.697 0.9491 2.522 1.343
Lee-SKKU-LG 6 0.8079 17.449 3.612 0.9718 5.283 1.576
CLone 7 0.6339 46.486 8.899 0.9332 16.119 2.730
Rzjs 8 0.6787 48.742 7.963 0.9287 18.322 2.806
Zeng-SJTU 9 0.5423 46.457 17.757 0.6235 37.966 12.306
Zou-SZU 10 0.5944 86.892 18.090 0.6708 115.814 22.005
watsons 11 0.4940 83.746 19.186 0.6165 90.970 16.478
Baseline - 0.9003 6.967 2.178 0.9804 2.423 1.299
Inter-observer - 0.9848 1.317 0.306 0.9996 1.069 0.474
Ensemble - 0.9364 4.055 1.347 0.9819 2.396 1.403
Figure 4: Distribution of segmentation errors for the top five performing algorithms on the CT test set. The first row shows the fragment-wise results. The second rows shows the anatomical results.
In addition, we evaluate the performance of our previously proposed CT semantic segmentation method, FracSegNet, described in Sec. 2.1, as a retrospective baseline. Note that the results reported in the original study were based on a different test set, and the evaluation ground truth in that work did not separate secondary fragments as required in this challenge.
4.3.2 X-ray segmentation performance
@@
Example X-ray segmentation results are shown in Fig. 5. The performance metrics of each algorithm submitted to the final test are summarized in Table 6 and Fig. 6.
Table 6: Quantitative results for the final test phase of the X-ray task. HD95 and ASSD values are reported in pixel unit. The best scores for each metric are shown in bold.
Team Rank ↓ IoU-F ↑ HD95-F ↓ ASSD-F ↓ IoU-A ↑ HD95-A ↓ ASSD-A ↓
SMILE 1 0.7736 37.366 8.545 0.9238 13.299 2.222
MedIG 2 0.7640 40.931 9.043 0.9268 15.979 2.462
LME 3 0.7150 42.986 10.355 0.8760 21.474 4.049
luckyjing 4 0.7145 45.752 10.398 0.8125 46.060 8.877
Puvvula-NIT 5 0.5506 158.171 28.092 0.7599 150.135 23.778
Ensemble - 0.8022 30.092 6.378 0.7970 69.878 12.268
Figure 6: Distribution of segmentation errors for algorithms on the X-ray test set. The first row shows the fragment-wise results. The second rows shows the anatomical results.
Figure 7: Distribution of the IoU-F score of the five top performing teams in both tasks across cases. In the X-ray task, the scores for the 20 synthetics samples per original CT scan are averaged.
Fig. 7 presents the distribution of IoU-F across test cases for both tasks. Among them, case 19 (d in Fig. 3 and Fig. 5) emerged as an outlier, exhibiting the lowest segmentation accuracy in both tasks. This was mainly attributed to the presence of multiple small fragments, particularly in the ramus pubis region.
To assess the statistical significance of pairwise differences between competing teams, we employed the Wilcoxon signed-rank test on the IoU-F scores. We specifically conducted a one-sided Wilcoxon signed-rank test to determine whether the IoU-F scores of one team were statistically significantly higher than those of another. In this setting, the null hypothesis states that the median difference between paired observations is zero or negative. The alternative hypothesis asserts that the IoU-F scores of the first team are systematically greater than those of the second team. The resulted significance matrices are shown in Fig. 8. A p-value below 0.05 was considered statistically significant.
@@
4.4 Inter-observer agreement
In order to assess the quality of the ground truth, as well as to compare the algorithm’s performance against human experts’, we perform a retrospective analysis to assess inter-observer variability by manually annotating the CT test set with an additional independent annotator, using the same protocol described in Sec. 3.2.2.
The assessment results are also summarized in Table 5. The high agreement between annotators confirmed the reliability of the ground truth. The primary source of variation stemmed from differences in determining whether an incomplete fracture should be labeled as two separate fragments or a single fragment in certain cases.
@@ 4.4 Inter-observer agreement
4.5 Impact of fragment size and count on segmentation
@@
Figure 9: Impact of fragment size and count on the IoU-F metric for the top five performing algorithms in both tasks. In (a) and (c), fragment sizes are binned in intervals of 500,000 mm3 for CT and 10,000 pixel3 for X-ray to enhance visual clarity.
4.6 False-positive analysis
Table 7: False-positive predictions in the final test phase. The mean and standard deviation of unmatched fragment counts are reported. The upper and lower sections show results for the CT and X-ray tasks, respectively.

### v2

@@ IV-B1 CT Segmentation
The authors formulated the task as two semantic segmentation problems. The first network performed anatomical segmentation of the three bone regions. The second network predicted the adaptive border boundary core (ABBC) representation, which consists of four semantic classes: background, boundary, border, and core. The core and boundary classes together define individual fragments, with the boundary class acting as a separator between the core and the background or adjacent fragments. The authors introduced a novel dynamic boundary thickness adjustment using the medial axis transform, ensuring that thin fragments remain connected within the core region. Given a prediction, connected component analysis was applied to the core class to extract initial fragment instances. The boundary class was then merged with the nearest seed. To ensure anatomical consistency, fragment instances were adjusted by splitting them along anatomical borders and merging smaller fragments with the closest anatomically correct instance. Additionally, the border class was used to refine the segmentation by modifying fragments through split-merge operations, ensuring alignment with the network predictions. [^5] MIC-DKFZ CT algorithm: https://codebase.helmholtz.cloud/hi-dkfz/applied-computer-vision-lab/challenges/abbc.
@@
The authors proposed a two-stage segmentation framework. The first stage focused on localizing the three bone regions. To achieve this, the authors trained a U-Net model on downsampled CT scans, which reduced the computational costs and provided a global context that prevented excessive false positives when processing smaller patches. The localized regions were then refined at full resolution for accurate anatomical segmentation. The second stage focused on fracture segmentation within each of the three bone regions. The authors categorized bone fragments into primary and secondary components. After obtaining initial predictions, connected component analysis was applied to reassign secondary fragment labels in accordance with evaluation rules. The final segmentation result was generated by combining the outputs from both stages. A final refinement step ensured that the fracture segmentation was aligned with the foreground mask from the first stage. [^6] SMILE CT algorithm: https://github.com/yuepeiyan/PENGWIN_Challenge.
@@
The authors employed a two-stage segmentation pipeline to accurately localize and segment fractured bone fragments. In the first stage, the TotalSegmentator [wasserthal2023totalsegmentator] model was used to localize the pelvic region, after which the three bone regions were cropped from the original CT scans. In the second stage, the NexToU model was applied to segment primary and secondary fragments within each cropped bone region. The largest connected component was extracted as the primary fragment, and the remaining secondary fragments were expected to be naturally separated. A connected component analysis step was used to isolate individual secondary fragments. [^7] MedIG CT algorithm: https://github.com/pzhhhhh2263/MICCAI-Challenge-PENGWIN2024.
@@
The authors developed a refinement-based segmentation pipeline, integrating anatomical segmentation with fracture surface segmentation to ensure anatomically consistent fracture identification. The approach utilized two deep networks. The first network, adapted from the FracSegNet study [liu2023pelvic], was used for segmenting the three bone regions. The second network, trained on fracture surfaces extracted from the PENGWIN dataset, identified fracture boundaries. For training, the authors generated a complete anatomical segmentation of the pelvis by merging all bone fragments. They then extracted contours of both individual fragments and the anatomical segmentation, removing overlapping edges to isolate fracture surfaces—where fractured bones were in contact. The fracture network was trained to predict these fracture surfaces from CT scans. In post-processing, segmentations from both networks were merged. A continuity analysis was then performed to separate individual fractures, with fragments labeled according to their anatomical overlap. [^8] MedApp-AGH CT algorithm: https://github.com/Jarartur/pengwin-challenge-submission.
@@
All five teams employed a two-stage approach, first performing anatomical segmentation before segmenting fragments. As shown in Fig. 4, the key variations among their methods lay in the fracture segmentation step—whether to extract individual fragments directly or to first detect border regions and infer separation. In Table II, we summarize some of the main components of the algorithms, including network architecture, pre-processing, and training schemes. A detailed discussion of methodology is included in Sec. V-A1.
IV-B2 X-Ray Segmentation
Figure 5: Different fracture segmentation pipelines in the X-ray task. (a) Bone extraction followed by primary-secondary segmentation by SMILE. (b) Direct semantic segmentation for the three bones in parallel by MedIG. (c) Bone extraction followed by instance segmentation by LME. (d) Direct instance segmentation with anatomical classification by luckyjing.
TABLE III: Summary of the top-performing participating algorithms for X-ray segmentation. Spatial augmentation includes random scaling, mirroring, and flipping. Intensity augmentation includes noising, blurring, and histogram adjustments.
@@ IV-B2 X-Ray Segmentation
The authors addressed the limitations of semantic segmentation in handling overlapping anatomical structures in X-rays by developing a two-stage hierarchical strategy similar to their CT solution. In the first stage, three dedicated U-Net models were trained to segment the three bone regions separately, allowing for focused substructure analysis within each area. In the second stage, a hierarchical fragment classification approach was implemented, adapted from the authors’ CT segmentation method, with modifications specific to X-ray imaging. Separate models were employed to distinguish primary and secondary fracture fragments, mitigating the overlapping structures in X-rays. While this approach effectively reduced primary-secondary overlap, challenges remained in cases where multiple secondary fragments overlapped. For post-processing, connected component analysis was applied to relabel secondary fragments, while primary fragment predictions were refined through consistency checks with the localization masks from the first stage, ensuring that only validated foreground pixels were retained. [^9] SMILE X-ray algorithm: https://github.com/yuepeiyan/PENGWIN_Challenge.
@@
The authors employed a multi-model approach to process each of the three bone regions in X-rays. Each model generated a multi-channel 3D tensor as output. The first channel represents the primary fragment, while the remaining nine channels corresponds to secondary fragments. The segmentation results from the three models were then combined to construct the final output. To optimize model performance, the authors applied as loss function an unweighted combination of Dice loss, binary CE loss, and focal Tversky loss. A pre-trained ImageNet model was used to initialize the network, enhancing feature extraction capabilities. The network architecture was based on a UNet++ backbone with a RegNetY-160 encoder, which provided a balance between spatial resolution and computational efficiency. [^10] MedIG X-ray algorithm: https://github.com/pzhhhhh2263/MICCAI-Challenge-PENGWIN2024.
@@
The authors proposed a category and fragment segmentation (CFS) framework. First, three separate networks, each based on the Swin-UNETR architecture, were used to segment the three bone regions. Then, a fragment segmentation network was applied to isolate the bone fragments within each region. The fragment segmentation step employed three separate Mask R-CNN-based networks, which performed both segmentation and object detection through bounding box prediction and classification. These networks took both the image and the binary anatomical mask as input, allowing the model to focus on specific bone regions while reducing interference from surrounding structures. During inference, the network predicted multiple bounding boxes, each associated with a confidence score. Only bounding boxes with a confidence score above 0.8 were retained for further analysis. In post-processing, the fragment segmentation masks were multiplied by the anatomical masks from the first step to ensure spatial consistency.
@@
The authors employed Mask R-CNN with a ResNet50 backbone for instance segmentation. The model consists of two main stages: a region proposal network that generates candidate object regions, and a detection network that refines these regions while performing anatomical classification, bounding-box regression, and mask prediction. Mask R-CNN was employed to predict pixel-wise binary masks for each region of interest (RoI). To maintain spatial alignment between RoIs and extracted features, the network replaced RoI Pooling with RoI Align, which eliminates quantization errors by using bilinear interpolation instead of discrete binning. This significantly improved mask prediction accuracy, particularly for small objects. To address the specific challenges of pelvic fragment segmentation, the loss weights between classification and segmentation were carefully adjusted during training. Additionally, a structured loss function combining IoU and cross-entropy loss was incorporated, enhancing the model’s robustness and segmentation quality in occluded scenarios.
@@
As shown in Fig. 5, the solutions in the X-ray task exhibited greater diversity in their overall approach than the CT task. The main components of the algorithms are summarized in Table III and further discussed in Sec. V-A2.
IV-C Performance Analysis
IV-C1 CT Segmentation Performance
Figure 6: Example CT segmentation results from the top-performing teams. Four representative cases are shown in (a)-(d), illustrating varying levels of complexity from easy to difficult. A complex fractured ramus pubicus region is shown in (e).
Example CT segmentation results are shown in Fig. 6. The performance metrics of each algorithm submitted to the final test are summarized in Table IV. In addition, we evaluate the performance of our previously proposed CT semantic segmentation method, FracSegNet, described in Sec. II-A, as a retrospective baseline.
TABLE IV: Quantitative results for the final test phase of the CT task. HD95 and ASSD values are reported in millimeters. Container runtime is reported in seconds. The best scores for each metric are shown in bold.
Team Rank ↓ IoU-F ↑ HD95-F ↓ ASSD-F ↓ IoU-A ↑ HD95-A ↓ ASSD-A ↓ Time ↓
MIC-DKFZ 1 0.9296 5.866 1.843 0.9810 2.368 1.290 510.4
SMILE 2 0.9096 5.562 1.663 0.9802 2.410 1.321 333.1
MedIG 3 0.9084 6.207 1.680 0.9798 2.452 1.298 257.4
MedApp-AGH 4 0.9049 6.899 2.007 0.9783 2.398 1.385 617.2
Sano 5 0.8802 6.866 1.697 0.9491 2.522 1.343 308.1
Lee-SKKU-LG 6 0.8079 17.449 3.612 0.9718 5.283 1.576 535.5
CLone 7 0.6339 46.486 8.899 0.9332 16.119 2.730 160.0
Rzjs 8 0.6787 48.742 7.963 0.9287 18.322 2.806 343.5
Zeng-SJTU 9 0.5423 46.457 17.757 0.6235 37.966 12.306 206.9
Zou-SZU 10 0.5944 86.892 18.090 0.6708 115.814 22.005 360.7
watsons 11 0.4940 83.746 19.186 0.6165 90.970 16.478 255.4
Baseline - 0.9003 6.967 2.178 0.9804 2.423 1.299 -
Inter-observer - 0.9848 1.317 0.306 0.9996 1.069 0.474 -
IV-C2 X-Ray Segmentation Performance
@@
Example X-ray segmentation results are shown in Fig. 7. The performance metrics of each algorithm submitted to the final test are summarized in Table V. To assess the statistical significance of pairwise differences between competing teams, we conducted a one-sided Wilcoxon signed-rank test to determine whether the IoU-F scores of one team were statistically significantly higher than those of another. The significance matrices are shown in Fig. 8. A p-value below 0.05 was considered statistically significant.
TABLE V: Quantitative results for the final test phase of the X-ray task. HD95 and ASSD values are reported in pixel unit. Container runtime is reported in seconds. The best scores for each metric are shown in bold.
Team Rank ↓ IoU-F ↑ HD95-F ↓ ASSD-F ↓ IoU-A ↑ HD95-A ↓ ASSD-A ↓ Time ↓
SMILE 1 0.7736 37.366 8.545 0.9238 13.299 2.222 207.4
MedIG 2 0.7640 40.931 9.043 0.9268 15.979 2.462 168.5
LME 3 0.7150 42.986 10.355 0.8760 21.474 4.049 210.9
luckyjing 4 0.7145 45.752 10.398 0.8125 46.060 8.877 141.3
Puvvula-NIT 5 0.5506 158.171 28.092 0.7599 150.135 23.778 140.8
@@
IV-D Inter-Observer Agreement
In order to assess the quality of the ground truth, as well as to compare the algorithm’s performance against human experts’, we perform a retrospective analysis to assess inter-observer variability by manually annotating the CT test set with an additional independent annotator, using the same protocol described in Sec. III-B2.
The assessment results are also summarized in Table IV. The high agreement between annotators confirmed the reliability of the ground truth. The primary source of variation stemmed from differences in determining whether an incomplete fracture should be labeled as two separate fragments or a single fragment in certain cases.
@@ IV-D Inter-Observer Agreement
IV-E Impact of Fragment Size and Count on Segmentation
@@
Figure 9: Impact of fragment size and count on the IoU-F metric for the top five performing algorithms in both tasks. In (a) and (c), fragment sizes are binned in intervals of 500,000 mm3 for CT and 10,000 pixel2 for X-ray to enhance visual clarity.
IV-F False-Positive Analysis
TABLE VI: False-positive predictions in the final test phase. The mean and standard deviation of unmatched fragment counts are reported. The upper and lower sections show results for the CT and X-ray tasks, respectively.